# A Universal Scheme for Dynamic Partitioned Shortest Path Index

### v2

@@ 2.3. Traditional PSP Strategy
Note that the construction of $\tilde{L}$ (Step 3, Step 4) can be parallelized with $L_{i}$ (Step 2) since they are independent and both rely on Step 1. All pairs of boundaries are independent of each other so we only need $|B|$ times of Dijkstra’s with time complexity $O(nlogn+m)$. Then each partition’s label $L_{i}$ can be constructed in parallel, and $\tilde{L}$’s construction is also independent to them, so its complexity is the worst case of them: $max\{O_{c}(G_{i}),O_{c}(\tilde{G})\}$, where $O_{c}$ is the complexity of underlying index’s construction time as our discussion is not fixed to any specific index type.
@@
Query Processing. By referring to the cut property of boundary vertex, the shortest path queries are divided into two categories and processed by utilizing $L$:
Case 1: Same-Partition i.e., $\forall s,t\in G_{i}$, $Q(s,t)=d_{L_{i}}(s,t)$;
Case 2: Cross-Partition, i.e., $\forall s\in G_{i},t\in G_{j}(i\neq j),Q(s,t)=$
$\begin{cases}d_{\tilde{L}}(s,t)&s,t\in B\\ \min\limits_{b_{q}\in B_{j}}\{d_{\tilde{L}}(s,b_{q})+d_{L_{j}}(b_{q},t)\}&s\in B,t\notin B\\ \min\limits_{b_{p}\in B_{i}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},t)\}&s\notin B,t\in B\\ \min\limits_{b_{p}\in B_{i},b_{q}\in B_{j}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},b_{q})+d_{L_{j}}(b_{q},t)\}&s\notin B,t\notin B\\ \end{cases}$
@@
In summary, when $s$ and $t$ are in the same partition, we can use $L_{i}$ to answer $d_{G}(s,t)$; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The same-partition query complexity is $O_{q}(L_{i})$, where $O_{q}$ is the index’s query complexity. The cross-partition query is made up of three parallel query sets, and the complexity is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. The proof of the query processing can be found in the Appendix of our extended version (Zhang et al., 2023b).
2.4. Shortest Path Algorithms
We summarize the shortest path algorithms as follows: 1) Direct Search such as Dijkstra’s (Dijkstra, 1959) and $A^{*}$ (Hart et al., 1968) searches the graph when the index cannot be constructed due to the large graph size (QbS (Wang et al., 2021), ParDist (Chondrogiannis and Gamper, 2016)) or complicated problem (COLA (Wang et al., 2016)). It takes no time in index update, but queries are slowest; 2) Contraction Hierarchy (CH) (Geisberger et al., 2008) is a widely used lightweight index that contracts the vertices in a pre-defined order and preserves the shortest path information by adding shortcuts among the contracted vertex’s neighbors. The search-based CH (Geisberger et al., 2008) has a small index size but takes longer to build and maintain (Zhang et al., 2021a; Geisberger et al., 2012) while concatenation-based CH (Ouyang et al., 2020; Wei et al., 2020) is much faster to build and maintain if the tree-width is small. CH’s query performance is generally around 10$\times$ faster than the direct search but much slower than the hub labelings. Surprisingly, no PSP index has used CH as underlying index; 3) Pruned Landmark Labeling (PLL). Although many hub labeling methods have been proposed in the past decade, only two are widely used. Built by either pruned search (Akiba et al., 2013) or propagation (Li et al., 2019b; Jin et al., 2020; Zhang et al., 2021b, 2023a), PLL is the only index that can work on the graph with large treewidth. Therefore, graphs with this property use it as the underlying index (e.g., QsB (Wang et al., 2021; Farhan et al., 2022) for its landmarks, Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) for its core); 4) Tree Decomposition (TD) (Ouyang et al., 2018; Chang et al., 2012; Wei, 2012; Chen et al., 2021; Zhang et al., 2021c). As another widely used hop labeling, it is much faster than PLL for graphs with smaller treewidth but cannot scale to large ones. For instance, CT (Li et al., 2020a; Zheng et al., 2022) uses it for its periphery, and FHL (Liu et al., 2021, 2022a) use it as a forest; 5) All Pair Pre-computation pre-computes all-pair distance with the fastest query, but incurs long index construction and large index storage. G-Tree (Zhong et al., 2013, 2015) and ROAD (Lee et al., 2009, 2010) store the distance between boundary vertices in each partition.
3. Novel PSP Strategies
As stated in Section 2, the traditional PSP strategy can be time-consuming for large networks since it requires a laborious computation of all-pair boundary distances using Dijkstra’s search. To tackle this problem, we propose two novel PSP strategies and analyze how they could be leveraged to correctly answer the shortest path queries. Besides, we design the index update approaches systematically to support the dynamic application of both the existing and novel PSP strategies. Lastly, we proposed a pruning technique to slim the overlay graph to achieve better index performance. It should be noted that the PSP strategy is a general procedure that any PSP index needs to follow and can work with any type of graph partition method or SP algorithm. So we introduce the PSP strategies without mentioning the specific partitioning and SP approach.
3.1. No-Boundary and Post-Boundary Strategy
@@
Since the traditional partitioned index starts by precomputing the all-pair distance among boundary vertices, we call this approach Pre-Boundary Strategy. It seems that this strategy, with the theoretical guarantee for correct partitioned shortest-path processing, could handle all the partitioned queries well. Nevertheless, the first step of the Pre-Boundary could be very time-consuming because only the index-free SP algorithms like Dijkstra’s (Dijkstra, 1959) or A (Hart et al., 1968) could be utilized. As a result, the index construction and maintenance efficiency suffer when the graph has numerous boundary vertex pairs. However, it appears that pre-computing the all-pair boundary distance is an essential procedure for constructing the “correct” PSP index, otherwise the correctness of the subgraph index $L_{i}$ cannot be guaranteed because the shortest distance between boundary vertices within one partition could pass through another partition. For instance, as shown in Figure 3, $p(v_{6},v_{11})=\left<v_{6},v_{12},v_{11}\right>$ goes outside $G_{4}$ and passes through $v_{12}\in G_{3}$. As a result, $d(v_{6},v_{11})$ cannot be answered correctly only with $G_{3}$ and the global shortest distance $d(v_{6},v_{11})$ calculation is essential before constructing $L_{3}$. Then a question naturally arises: do we really require precomputing the boundary all-pair distance first? In other words, does there exist a chance to keep index correctness by dropping this time-consuming step? Based on this brainstorm, we break the traditional misconception and propose a novel No-Boundary Strategy as follows, which significantly reduces the index construction/maintenance time by skipping the time-consuming pre-computation step.
@@
Index Construction. It contains three steps as shown in Figure 4. With a graph partitioned into subgraphs $\{G_{i}\}$, we start by constructing the shortest path index $L_{i}$ for each subgraph $G_{i}$ parallelly in Step 1; Then Step 2 construct the overlay graph $\tilde{G}$ based on $\{L_{i}\}$. For instance, $\tilde{G}$ is made up of boundary all-pair edges $(v_{3},v_{10})$, $(v_{8},v_{9})$, $(v_{5},v_{12})$, $(v_{6},v_{11})$ (with their weight calculated based on $L_{1},L_{2},L_{3}$, and $L_{4}$) and inter-edges. It should be noted that although the overlay graph $\tilde{G}$ has the same structure as that in the Pre-boundary Strategy, their edge weights differ as $\{L_{i}\}$ of them are constructed from different subgraphs; Step 3 constructs the shortest path index $\tilde{L}$ for $\tilde{G}$. Since the partition indexes $\{L_{i}\}$ are constructed in parallel first followed by the construction of $\tilde{L}$, the No-Boundary takes $max\{O_{c}(G_{i})\}+O_{c}(\tilde{G})$ time in index construction.
@@
Query Processing. Now that $L_{i}$ cannot answer $G_{i}$’s query correctly as the global distance between boundary all-pair is not covered, then how can No-Boundary answer query correctly? Before revealing it, we first prove that the correctness of $\tilde{L}$ still holds even though it is built upon the incorrect $\{L_{i}\}$, that is the queries between boundary vertices can be correctly processed with $\tilde{L}$, using:
@@
We divide all the scenarios into three cases as shown in Figure 5-(a): 1) $s,t\in G_{i}$ and $p_{G}(s,t)$ only passes through the interior of $G_{i}$, then it is obvious that $d_{G}(s,t)=d_{G_{i}}(s,t)$ holds. Since $\tilde{w}(s,t)=d_{G_{i}}(s,t)$ is leveraged for $\tilde{G}$’s construction, we can obtain that $d_{G}(s,t)=d_{\tilde{G}}(s,t)=d_{\tilde{L}}(s,t)$; 2) $s,t\in G_{i}$ with $p_{G}(s,t)$ going outside of $G_{i}$; 3) $s\in G_{i}$ and $t\in G_{j}(i\neq j)$. In the latter two cases, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$). For two adjacent vertices $b_{i},b_{j}\in p_{c}$, if $b_{i}$ and $b_{j}$ are in the same partition, then $\tilde{w}(b_{i},b_{j})$ can be correctly obtained in $\tilde{G}$ via case 1); otherwise, $(b_{i},b_{j})$ is an inter-edge with $\tilde{w}(b_{i},b_{j})=$$w(b_{i},b_{j})$ naturally correct. Thus, the shortest distance among boundary vertices can be correctly calculated by accumulating the edge weights on $\tilde{G}$. ∎
@@
Based on the above Theorem, we continue to answer the shortest distance queries by discussing all scenarios in two cases:
Case 1: Same Partition $s,t\in G_{i}$, we report $Q(s,t)$ as follows:
@@
We denote $d_{L_{i}}(s,t)$ as $d_{2}$, $\min\{d_{L_{i}}(s,b_{i1})+d_{\tilde{L}}(b_{i1},b_{i2})+d_{L_{i}}(b_{i2},t)\}$ as $d_{4}$, and divide all the scenarios into two subcases:
Subcase 1: $p_{G}(s,t)$ does not go outside of $G_{i}$, as shown in the first row of Figure 5. No matter $s$ and $t$ are borders or not, $d_{L_{i}}(s,t)$ (i.e., $d_{2}$) is enough to answer $d_{G}(s,t)$ as $L_{i}$ is built based on $G_{i}$ which contains all necessary information for finding the shortest path.
@@
Case 2: Cross-Partitions, i.e., $s\in G_{i},t\in G_{j},i\neq j$, we process queries the same manner as that of Pre-boundary Strategy.
@@
Note that the proof of Lemma 3 and all the other Lemmas of this paper are available in the Appendix of our extended version (Zhang et al., 2023b). We can use $\tilde{L}$ to answer $d_{G}(s,t)$ when $s$ and $t$ are both boundary vertices; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The intra-query complexity is $max\{O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$, while the inter-query takes $O_{q}(\tilde{L})$ time when $s$ and $t$ are both boundary vertices, and the complexity of other case is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. Therefore, we make a bold attempt to enhance the performance of PSP index by skipping the heavy pre-computation, such that the index construction time is largely reduced so as the index maintenance (as will introduced in Section 3.2). Surprisingly, with incorrect distance value in the No-boundary index, we prove theoretically that it can support the query answering with correctness guarantee.
@@
By comparing the query processing procedure of the traditional Pre-boundary strategy and our proposed No-boundary strategy, we find that it is more complex for No-boundary to process those queries with endpoints in the same partition because of the incorrectness of $\{L_{i}\}$. We repair this weakness by fixing the incorrect boundary all-pair distance in each partition $G_{i}$ and then transforming $\{L_{i}\}$ to their correct version. Specifically, we calculate the boundary all-pair distance $d(b_{i1},b_{i2})$ within each partition $G_{i}$ by leveraging $\tilde{L}$ and insert its corresponding edge $e(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$. For instance, $d(v_{3},v_{10})$, $d(v_{8},v_{9})$, $d(v_{5},v_{12})$ and $d(v_{6},v_{11})$ are calculated based on $\tilde{L}$ and these edges (as colored red in Figure 4) are inserted into their corresponding partitions with new subgraphs $\{G^{\prime}_{i}\}$ formed. Followed by, we treat those newly inserted edges as graph updates and refresh $\{L_{i}\}$ to $\{L^{\prime}_{i}\}$ by invoking the index maintenance algorithms (Ouyang et al., 2020; Zhang et al., 2021c; Zhang and Yu, 2022; Farhan et al., 2022) in each partition parallelly. As these procedures are implemented after the No-boundary, we call this strategy as Post-boundary. By referring to Theorem 1, the correct global distance information is contained in $\{G^{\prime}_{i}\}$, so $\{L^{\prime}_{i}\}$ can support the correct query processing of $Q(s,t)$ with $s,t\in G_{i}$. With the consistent query procedure as that of Pre-boundary, the query efficiency of Post-boundary reaches the state-of-the-art level.

### v3

@@ 5.1. Traditional Pre-Boundary Strategy
Note that the construction of $\tilde{L}$ (Step 3, Step 4) can be parallelized with $L_{i}$ (Step 2) since they are independent and both rely on Step 1. For Step 1, we need $|B|$ times of Dijkstra’s with time complexity $O(|B|\cdot(n\log n+m))$. Then each partition’s label $L_{i}$ can be constructed in parallel, and $\tilde{L}$’s construction is also independent of them, so its complexity is the worst case of them: $max\{O_{c}(G_{i}),O_{c}(\tilde{G})\}$, where $O_{c}$ is the complexity of underlying index’s construction time as our discussion is not fixed to any specific index type.
@@
Query Processing. The shortest path queries can be divided into two categories and processed by utilizing $L$:
[…]
@@
In summary, when $s$ and $t$ are in the same partition, we can use $L_{i}$ to answer $d_{G}(s,t)$; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The same-partition query complexity is $O_{q}(L_{i})$, where $O_{q}$ is the index’s query complexity. The cross-partition query is made up of three parallel query sets, and the complexity is the worst of them: $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$. The proof of the query processing is in the Appendix of our extended version (Zhang et al., [n. d.]). Since the traditional partitioned index starts by precomputing the all-pair distance among boundary vertices, we call this approach Pre-Boundary Strategy.
5.2. Novel No-Boundary and Post-Boundary Strategies
@@
The pre-boundary strategy with the theoretical guarantee for correct partitioned shortest-path processing could handle all the queries well. However, the Step 1 of the Pre-Boundary could be very time-consuming because only the index-free SP algorithms like Dijkstra’s (Dijkstra, 1959) or A (Hart et al., 1968) could be utilized. As a result, the index construction and maintenance efficiency suffer when the graph has numerous boundary vertex pairs. However, it appears that pre-computing the all-pair boundary distance is an essential procedure for constructing the “correct” PSP index. Otherwise, the correctness of the subgraph index $L_{i}$ cannot be guaranteed because the shortest distance between boundary vertices within one partition could pass through another partition. For instance, as shown in Figure 4, $sp(v_{6},v_{11})=\left<v_{6},v_{12},v_{11}\right>$ goes outside $G_{4}$ and passes through $v_{12}\in G_{3}$. As a result, $d(v_{6},v_{11})$ cannot be answered correctly only with $G_{3}$ and the global shortest distance $d(v_{6},v_{11})$ calculation is essential before constructing $L_{3}$. Then, a question arises naturally: do we really require precomputing the boundary all-pair distance first? In other words, is there a chance to keep index correctness by dropping this time-consuming step? Based on this brainstorm, we break the traditional misconception and propose a novel No-Boundary Strategy as follows, which significantly reduces the index construction/maintenance time by skipping the time-consuming pre-computation step.
@@
Index Construction. It contains three steps as shown in Figure 5. With a graph partitioned into subgraphs $\{G_{i}\}$, we start by constructing the partition index $L_{i}$ for each subgraph $G_{i}$ parallelly in Step 1; Then Step 2 construct the overlay graph $\tilde{G}$ based on $\{L_{i}\}$. For instance, $\tilde{G}$ is made up of boundary all-pair edges $(v_{3},v_{10})$, $(v_{8},v_{9})$, $(v_{5},v_{12})$, $(v_{6},v_{11})$ (with their weight calculated based on $L_{1},L_{2},L_{3}$, and $L_{4}$) and inter-edges. It should be noted that although the overlay graph $\tilde{G}$ has the same structure as that in the Pre-boundary Strategy, their edge weights differ as $\{L_{i}\}$ of them are constructed from different subgraphs; Step 3 constructs the overlay index $\tilde{L}$ for $\tilde{G}$. Since the partition indexes $\{L_{i}\}$ are constructed in parallel first followed by the construction of $\tilde{L}$, the No-Boundary takes $max\{O_{c}(G_{i})\}+O_{c}(\tilde{G})$ time in index construction.
@@
Query Processing. Now that $L_{i}$ cannot answer $G_{i}$’s query correctly as the global distance between boundary all-pair is not covered, then how can No-Boundary answer query correctly? Before revealing it, we first use the following theorem to prove that the correctness of $\tilde{L}$ still holds even though it is built upon the incorrect $\{L_{i}\}$, i.e., the queries between boundary vertices can be correctly processed with $\tilde{L}$.
@@
We divide all the scenarios into three cases as shown in Figure 6-(a): 1) $s$ and $t\in G_{i}$ and $p_{G}(s,t)$ only passes through the interior of $G_{i}$, then it is obvious that $d_{G}(s,t)=d_{G_{i}}(s,t)$ holds. Since $\tilde{w}(s,t)=d_{G_{i}}(s,t)$ is leveraged for $\tilde{G}$’s construction, we can obtain that $d_{G}(s,t)=d_{\tilde{G}}(s,t)=d_{\tilde{L}}(s,t)$; 2) $s$ and $t\in G_{i}$ with $p_{G}(s,t)$ going outside of $G_{i}$; 3) $s\in G_{i}$ and $t\in G_{j}(i\neq j)$. In the latter two cases, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$). For two adjacent vertices $b_{i},b_{j}\in p_{c}$, if $b_{i}$ and $b_{j}$ are in the same partition, then $\tilde{w}(b_{i},b_{j})$ can be correctly obtained in $\tilde{G}$ via case 1); otherwise, $(b_{i},b_{j})$ is an inter-edge with $\tilde{w}(b_{i},b_{j})=w(b_{i},b_{j})$ naturally correct. Thus, the shortest distance among boundary vertices can be correctly calculated by accumulating the edge weights on $\tilde{G}$. ∎
@@
Based on the above Theorem, we answer the shortest distance queries by discussing all scenarios in two cases:
Case 1: Same Partition $s,t\in G_{i}$, we report $Q(s,t)$ by distance concatenation as shown in the following theorem:
@@
We denote $d_{L_{i}}(s,t)$ as $d_{2}$, $\min\{d_{L_{i}}(s,b_{i1})+d_{\tilde{L}}(b_{i1},b_{i2})+d_{L_{i}}(b_{i2},t)\}$ as $d_{4}$, and consider two subcases:
Subcase 1: $p_{G}(s,t)$ does not go outside of $G_{i}$, as shown in the first row of Figure 6. No matter $s$ and $t$ are borders or not, $d_{L_{i}}(s,t)$ (i.e., $d_{2}$) is enough to answer $d_{G}(s,t)$ as $L_{i}$ is built based on $G_{i}$ which contains all necessary information for finding the shortest path;
@@
Case 2: Cross-Partitions, i.e., $s\in G_{i},t\in G_{j},i\neq j$, we process them in the same manner as Pre-Boundary Strategy.
@@
Proof.
We prove and discuss it with three subcases as shown in Figure 6-(c).
Subcase 1: $s,t\in B$, then $d_{G}(s,t)$ can be correctly answered according to Theorem 1;
Subcase 2: $s\in B$ or $t\in B$. As shown in the second illustration in Figure 6-(c), suppose $s$ is an inner vertex of $G_{i}$ and $t$ is a boundary vertex, we take the concise $p_{s,t}$ by extracting the boundary vertices as $p_{c}=\left<s,b_{0},\dots,b_{n},t\right>$. Then $b_{0}\in B_{i}$ and $p_{s,t}$ can be treated as concatenated by two sub-paths $p_{s,b_{0}}\oplus p_{b_{0},t}$. Specifically, $d_{G}(s,b_{0})=d_{L_{i}}(s,b_{0})$ by referring to the Subcase 1 of Case 1, and $d_{G}(b_{0},t)=d_{\tilde{L}}(b_{0},t)$ by referring to Theorem 1;
Subcase 3: $s\notin B,t\notin B$, which is the extended case of Subcase 2, so $d_{G}(s,t)$ can also be correctly answered. ∎
@@
For cross-partition queries, we can use $\tilde{L}$ to answer $d_{G}(s,t)$ when $s$ and $t$ are both boundary vertices; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The intra-query complexity is $max\{O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$, while the inter-query takes $O_{q}(\tilde{L})$ time when $s$ and $t$ are both boundary vertices, and the complexity of other case is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. Therefore, we make a bold attempt to enhance the performance of PSP index by skipping the heavy pre-computation, such that the index construction time is largely reduced so as the index maintenance (as will introduced in Section 5.3). Surprisingly, with incorrect distance value in the No-boundary index, we prove theoretically that it can support the query answering with correctness guarantee.
@@
By comparing the query processing procedure of the traditional Pre-boundary strategy and our proposed No-boundary strategy, we find that it is more complex for No-boundary to process those queries with endpoints in the same partition because of the incorrectness of $\{L_{i}\}$. We repair this weakness by fixing the incorrect boundary all-pair distance in each partition $G_{i}$ and then transforming $\{L_{i}\}$ to their correct version. Specifically, we calculate the boundary all-pair distance $d(b_{i1},b_{i2})$ within each partition $G_{i}$ by leveraging $\tilde{L}$ and insert its corresponding edge $e(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$ (Step 4). For instance, $d(v_{3},v_{10})$, $d(v_{8},v_{9})$, $d(v_{5},v_{12})$ and $d(v_{6},v_{11})$ are calculated based on $\tilde{L}$ and these edges (as colored red in Figure 5) are inserted into their corresponding partitions with new subgraphs $\{G^{\prime}_{i}\}$ formed. Followed by, we treat those newly inserted edges as graph updates and refresh $\{L_{i}\}$ to $\{L^{\prime}_{i}\}$ by invoking the index maintenance algorithms (Ouyang et al., 2020; Zhang et al., 2021c; Zhang and Yu, 2022; Farhan et al., 2022; Zhang et al., 2021b; Zhou et al., 2024b) in each partition parallelly (Step 5). As these procedures are implemented after the No-boundary, we call this strategy as Post-boundary. By referring to Theorem 1, the correct global distance information is contained in $\{G^{\prime}_{i}\}$, so $\{L^{\prime}_{i}\}$ can support the correct query processing of $Q(s,t)$ with $s,t\in G_{i}$. With the consistent query procedure as that of Pre-boundary, the query efficiency of Post-boundary reaches the state-of-the-art level.